OpenStudy (anonymous):

The area of a rectangle must be at least 135 square feet. If the length is 9 feet, find the minimum for the rectangle's width

OpenStudy (anonymous):

area of triangle=1/2 length *base 135=9*base*1/2

OpenStudy (neo92):

area of a rectangl = length*width here lenfth = 9 = constant therefore, area directly proportional to width (area)min = 9*(width)min 135 = 9*width width = 135/9 = 15 ft

OpenStudy (anonymous):

thanks

OpenStudy (neo92):

:D